OpenStudy (anonymous):

I need help finding the domain of the rational function 5x-7/4. Please! Thanks! :)

OpenStudy (anonymous):

\[\frac{5x-7}{4}=\frac{5}{4}x-\frac{7}{4}\] is not a rational function, it is a line

OpenStudy (anonymous):

I think it will be \[(-\infty , \infty )\] cuz no matter what numbers you blog the denominator will remain 4 it will never be zero and the rational function domain is real numbers except what makes the denominator zero .... right guys ?